# A ripple in time: a discontinuity in American history

### v2

@@
A ripple in time: a discontinuity in American history
Alexander Kolpakov Einsof Computing Inc., Toronto ON, Canada  and  Igor Rivin Department of Mathematics, Temple University, USA
(Date: January 19, 2024)
We would like to thank Google Colab for the opportunity to perform GPU computations, and HuggingFace for the opportunity to access their collection of pre–trained NLP models.
In this note we use the State of the Union Address (SOTU) dataset from Kaggle to make some surprising (and some not so surprising) observations pertaining to the general timeline of American history, and the character and nature of the addresses themselves. Our main approach is using vector embeddings, such as BERT (DistilBERT) [2, 7] and GPT–2 [6].
@@
It is worth noting that SOTU addresses provide relatively small writing samples (with about 8000 words on average, and varying widely from under 2000 words to more than 20000), and that the amount of authors is relatively large (we used SOTU addresses of 42 US presidents). This shows that the techniques employed turn out to be rather efficient, while all the computations described in this note can be performed using a single GPU instance of Google Colab.
@@
1. Problems and methods
1.1. Dataset
The dataset is exactly what you might think: it is the collection of complete texts of State of the Union addresses (from now on, SOTU) of US presidents, from John Adams to Donald Trump [9]. We modified the dataset by splitting the addresses into train and test files, such that each President has at least one train SOTU, and at least one test SOTU file (we use the same file for training and testing for Taylor, as it seems that splitting it into, say, halves makes even less sense). Also, some presidents are more represented than others in the SOTU corpus because they spent more years in office. On average, we keep the train/test split about 75%/25%. Also, we keep this split for each individual President whenever possible.
1.2. Methods of study
@@
To complete this investigation, we fine-tuned the DistilBERT model to detect which President delivered which address. The methods and results are described in Section 3. We also used DistilBERT for embeddings instead of the actual BERT as it is much faster and lightweight. Thus, our experiments can be easily replicated in a Google Colab notebook within reasonable time.
2. Results and observations
2.1. Temporal clustering
The first thing we saw was largely unsurprising: Addresses delivered by the same person were close together, and addresses chronologically close together were also close together (where we measure ”closeness” by the cosine distance between the vectors). This seems a natural consequence of the fact that language does change over time.
What was surprising however, was that there is a large break, as demonstrated by the UMAP visualization (Figures 0(a)0(b)), between addresses written before 1927 and addresses written after 1932. We were rather curious to see this, and did cluster analysis on the original vector embedding (without reducing by UMAP), and the results were confirmed (with a couple of minor differences) – see Figure 2.
@@ 2.1. Temporal clustering
Figure 1. UMAP visualizations of the GPT–2 embedding of SOTU
@@
Figure 2. Temporal clustering of SOTU embeddings
2.2. BERT vs GPT–2
Even though BERT is widely assumed to be the right kind of embedding for NLP classification tasks, we can clearly see that GPT–2 in conjunction with UMAP produces very good clustering, too. However, the clustering in the initial DistilBERT embedding has a much cleaner temporal separation than in the case of GPT–2 (compare Figure 1(a) to Figure 1(b)).
@@ 2.2. BERT vs GPT–2
3. Authorship attribution of the addresses
@@
We should note here that the amount of training data for each author is relatively small, and also varies greatly: from 1790 words in an average SOTU address for John Adams to 22614 words on average for William H. Taft [5]. The average length of a SOTU is $\sim$ 8358 words.
For comparison, the study carried out in [8] used novels for authorship attribution. The average words length of a novel is believed to be over 40000. The amount of authors for authorship attribution in [8] is only 8, with classification accuracy of around 85%-90%. Here, we manage to identify 42 authors with at least 93% accuracy based on a much smaller amount of data than 6 novels written by each.
4. Year of writing the addresses
@@
5. Discussion
How do we explain our results? The short answer is that we do not know. The long answer is that we have a couple of theories.
The first is that before the time of Franklin Roosevelt, it was extremely uncommon for presidents to use speech writers, while Roosevelt used their help a lot. It seems plausible that a speech written by a hired hand will be different from one written by the President of the United States (even though the President obviously maintained considerable editorial control to the present day).
@@ 5. Discussion
The colors in, say, the UMAP embedding of SOTU (Figures 0(a) and 0(b)) changes continuously representing the language shift. The break is however sharp, and thus likely comes for a different reason than just the language of time.
@@ Appendix A Dimension reductions of GPT–2 embedding
Figure 3. 2D visualizations of the GPT–2 embedding of SOTU
The most important part is the temporal aspect of SOTU speeches, not their clustering per se. This is shown in Figure 5, where we see that the cleanest separation belongs to UMAP. However, TriMAP and PaCMAP are not much worse: in fact, the three cases turn out to be quite comparable on the average run.
@@
Figure 4. 3D visualizations of the GPT–2 embedding of SOTU
@@
Figure 5. Temporal clustering of SOTU speeches: GPT–2 embedding followed by a dimension reduction technique
@@ Appendix B Dimension reductions of DistilBERT embedding
Figure 6. 2D visualizations of the DistilBERT embedding of SOTU
@@
Figure 7. 3D visualizations of the DistilBERT embedding of SOTU
Again, we are interested in the temporal aspect of clustering that is illustrated in Figure 5. There the cleanest separation belongs to TriMAP. After making several runs, one can see that TriMAP largely preserves its supremacy over UMAP and PaCMAP.
@@
Figure 8. Temporal clustering of SOTU speeches: DistilBERT embedding followed by a dimension reduction technique
In all pictures, temporal separation happens at the break of 1920 and 1921, though only one of the pictures shows a clear watershed moment. For the original DistilBERT embedding it is not quite the case: the temporal separation in Figure 1(b) is quite clean, although it happens about 16 years after 1921.

### v5

@@
Abstract
In this technical note we suggest a novel approach to discover temporal (related and unrelated to language dilation) and personality (authorship attribution) in historical datasets. We exemplify our approach on the State of the Union speeches given by the past 42 US presidents: this dataset is known for its relatively small amount of data, and high variability of the amount and style of texts. Nevertheless we manage to achieve about 95% accuracy on the authorship attribution task, and pin down the date of writing to a single presidential term.
I Results and novelty
In the present work we use the State of the Union Address (SOTU) dataset from Kaggle to make some surprising (and some expected) observations pertaining to the general timeline of American history, and the character and nature of the addresses themselves.
Our main approach is using vector embeddings, such as BERT (DistilBERT) [3, 9] and GPT–2 [8] in conjunction with dimension reduction techniques such as UMAP [6], TriMAP [1] and PaCMAP [13].
@@ I Results and novelty
Previous attempts at the authorship attribution tasks used much larger datasets [11] for historical analysis, or did not reach better accuracy except on a smaller “sure–fire” subset of the test data [10].
Other related studies [2, 4] concentrate on fine–tuning LLMs to handle temporal aspects of language (e.g. linking the names of country leaders to a given historical period) without an attempt to discover any relationships in the training data. Even though certain observations can likely be made by prompting the model, the latent space of the transformer is not analysed.
It is worth noting that SOTU addresses provide relatively small writing samples (with about 8000 words on average, and varying widely from under 2000 words to more than 20000), and that the number of authors is relatively large (we used SOTU addresses of 42 US presidents). This shows that the techniques employed turn out to be rather efficient, while all the computations described in this note can be performed using a single GPU instance of Google Colab.
@@
II Problems and methods
II-A Dataset
The dataset is the collection of complete texts of State of the Union addresses (from now on, SOTU) of US presidents, from John Adams to Donald Trump [12]. We modified the dataset by splitting the addresses into train and test files, such that each President has at least one train SOTU, and at least one test SOTU file (we use the same file for training and testing for Taylor, as it seems that splitting it into, say, halves makes even less sense). Also, some presidents are more represented than others in the SOTU corpus because they spent more years in office. On average, we keep the train / test split about 75% / 25%. Also, we keep this split for each individual President whenever possible.
II-B Methods of study
@@
To complete this investigation, we fine-tuned the DistilBERT model to detect which President delivered which address. The methods and results are described in Section IV. We also used DistilBERT for embeddings instead of the actual BERT as it is much faster and lightweight. Thus, our experiments can be easily replicated in a Google Colab notebook within reasonable time.
III Results and observations
III-A Temporal clustering
The first thing we saw was largely unsurprising: addresses delivered by the same person were close together, and addresses chronologically close together were also geometrically close together (where we measure ”closeness” by the cosine distance between the vectors). This seems a natural consequence of the temporal language dilation.
What was surprising, however, was that there is a large break, as demonstrated by the UMAP visualization (Figures 1(a)1(b)), between the addresses written before 1927 and those written after 1932. We were rather curious to see this, and did cluster analysis on the original vector embedding (without reducing by UMAP), and the results were confirmed (with a couple of minor differences) – see Figure 2.
@@ III-A Temporal clustering
Figure 1: UMAP visualizations of the GPT–2 embedding of SOTU
@@
Figure 2: Temporal clustering of SOTU embeddings
III-B BERT vs GPT–2
Even though BERT is widely assumed to be the right kind of embedding for NLP classification tasks, we can clearly see that GPT–2 in conjunction with UMAP produces very good clustering, too. However, the clustering in the initial DistilBERT embedding has a much clearer temporal separation than in the case of GPT–2 (compare Figure 2(a) to Figure 2(b)).
@@ III-B BERT vs GPT–2
IV Authorship attribution of the addresses
@@
We should note here that the amount of training data for each author is relatively small, and also varies greatly: from 1790 words in an average SOTU address for John Adams to 22614 words on average for William H. Taft [7]. The average length of a SOTU address is just about 8358 words.
For comparison, the study carried out in [11] used novels for authorship attribution. The average words length of a novel is believed to be over 40000. The number of authors for authorship attribution in [11] is only 8, with classification accuracy of around 85%-90%.
Another, less recent study [10] that targets SOTU addresses specifically, produces a comparably low accuracy. By the author’s own admission, it has a higher success rate of 95.7% as applied to an artificially created category based off similarity distances. The actual accuracy of the approach described in [10] can be computed as $(45+66+79)/(45+71+108)\approx 84.8\%$ (“Authorship attribution”, page 1652).
In the present work, we manage to identify all 42 authors with at least 93% accuracy in an honest–to–goodness test run, and with training based on a much smaller amount of data than 6 novels written by each.
V Year of writing the addresses
@@
VI Discussion
We will attempt to explain our results in the historical context, and make several conjectures about the root causes.
The first one is that before the time of Franklin Roosevelt, it was extremely uncommon for presidents to use speech writers, while Roosevelt used their help a lot. It seems plausible that a speech written by a hired hand will be different from one written by the President of the United States (even though US presidents obviously maintained considerable editorial control to the present day).
@@ VI Discussion
The colors in, say, the UMAP embedding of SOTU (Figures 1(a) and 1(b)) change continuously representing the language shift. However, the break is sharp, and thus likely comes for a different reason than just language dilation in time.
@@
Data accessibility
All numerical experiments can be reproduced by accessing the associated GitHub repository [5].
Authors’ contributions
A.K. and I.R. conceived the mathematical models, interpreted the computational results, and wrote the paper. A.K. and I.R. implemented and performed most of the numerical experiments, and otherwise contributed in equal parts.
@@ Appendix A Dimension reductions of GPT–2 embedding
Figure 3: 2D visualizations of the GPT–2 embedding of SOTU
The most important part is the temporal aspect of SOTU speeches, not their clustering per se. This is shown in Figure 5, where we see that the clearest separation belongs to UMAP. However, TriMAP and PaCMAP are not much worse: in fact, the three cases turn out to be quite comparable on the average run.
@@
Figure 4: 3D visualizations of the GPT–2 embedding of SOTU
@@
Figure 5: Temporal clustering of SOTU speeches: GPT–2 embedding followed by a dimension reduction technique
@@ Appendix B Dimension reductions of DistilBERT embedding
Figure 6: 2D visualizations of the DistilBERT embedding of SOTU
@@
Figure 7: 3D visualizations of the DistilBERT embedding of SOTU
Again, we are interested in the temporal aspect of clustering that is illustrated in Figure 5. There the clearest separation belongs to TriMAP. After making several runs, one can see that TriMAP largely preserves its supremacy over UMAP and PaCMAP.
@@
Figure 8: Temporal clustering of SOTU speeches: DistilBERT embedding followed by a dimension reduction technique
In all pictures, temporal separation happens at the break of 1920 and 1921, though only one of the pictures shows a clear watershed moment. For the original DistilBERT embedding it is not quite the case: the temporal separation in Figure 2(b) is quite clear, although it happens about 16 years after 1921.